OpenStudy (anonymous):

linear algebra mapping Let F : R^3 → R^3 be a linear mapping . Determine the imaging matrix F of the vector (1, 0 , -1) depicted at ( 2 , 2, 1 ) , (0 , 1, 2 ) depicted in ( 1 , 2, 1 ) and ( 1, 1, 0) is mapped onto itself.

ganeshie8 (ganeshie8):

So we are given the linear transformation of 3 vectors. (1, 0, -1) ---> (2, 2, 1) (0, 1, 2) ---> (1, 2, 1) (1, 1, 0) ---> (1, 1, 0) Your goal is to find the corresponding transformation matrix

OpenStudy (anonymous):

and I do that by checking the conditions? \[T(a ^{->}+b^{->})= T(a ^{->})+T(b^{->})\]

OpenStudy (anonymous):

Nice link, but do I have to do all that to figure out F?

ganeshie8 (ganeshie8):

Do you have any other better method ?

ganeshie8 (ganeshie8):

Notice that we can write (1, 0, 0) as linear combination of the 3 given vectors : (1, 0, 0) = 2(1, 0, -1) + 1(0, 1, 2) - 1(1, 1, 0)

ganeshie8 (ganeshie8):

Then, T(1, 0, 0) = 2T(1, 0, -1) + T(0, 1, 2) - T(1, 1, 0) = 2(2, 2, 1) + (1, 2, 1) - (1, 1, 0) = (4, 5, 3)

ganeshie8 (ganeshie8):

similarly you may try finding T(0, 1, 0) and T(0, 0, 1)

OpenStudy (anonymous):

@ParthKohli okey so we say A is our trans Matrix \[A \left(\begin{matrix}1 \\ 0 \\ -1\end{matrix}\right) = \left(\begin{matrix}2 \\ 2 \\1 \end{matrix}\right)\]and \[A \left(\begin{matrix}0 \\ 1 \\ 2\end{matrix}\right) = \left(\begin{matrix}1 \\ 2 \\ 1\end{matrix}\right)\]and also \[A \left(\begin{matrix}1 \\ 1 \\ 0\end{matrix}\right) = \left(\begin{matrix}1 \\ 1 \\ 0\end{matrix}\right)\] all together we have \[A\left[\begin{matrix}1 & 0 &1\\ 0 & 1&1 \\ -1 & 2&0\end{matrix}\right] = \left[\begin{matrix}2 & 1 &1\\ 2 & 2&1 \\ 1 & 1&0\end{matrix}\right]\] multiply the left with the inverse to solve for A \[A= \left[\begin{matrix}2 & 1 &1\\ 2 & 2&1 \\ 1 & 1&0\end{matrix}\right]*\left[\begin{matrix}1 & 0 &1\\ 0 & 1&1 \\ 1 & 2&0\end{matrix}\right] ^{-1}\] solve the inverse and then do the multiplication \[A=\left[\begin{matrix} 4& -3 &2\\ 5 & -4&3 \\ 3 & -3&2\end{matrix}\right] \]

ganeshie8 (ganeshie8):

Nicely done!